OpenStudy (anonymous):

what is 4z^2 + 16z +15 over 2z^2-z-15? OK I've narrowed it down to four choices 2z+3/z+3, 2z-3/z-3, 2z+3/2-3, 2z-3/z+3

OpenStudy (anonymous):

i got 17 .. O_O

OpenStudy (anonymous):

4z^2+16z+15 / 2z^2-z-15 4/2 =2 the z^2 cancel out . 16-1 = 15 the z cancels out . and 2+15 = 17 ? :o

OpenStudy (anonymous):

yea i got that too at first but its not right 8*(

OpenStudy (anonymous):

did you try -18 ?

OpenStudy (anonymous):

yea, that's not right either 8(, but its one of the four answers above

OpenStudy (anonymous):

Aww :l .. im sorry i tried . Idkk how to solve that one .